# Observe Then Act: Asynchronous Active Vision-Action Model for Robotic Manipulation

### v1

@@ IV-B Viewpoint-centric Voxel Alignment
Viewpoint movement expands the image observation space, increasing the variance of interaction samples[3]. Specifically, images captured from a static viewpoint are primarily influenced by the scene state, following the mapping $f^{\text{static}}:\mathcal{S}\rightarrow\mathcal{O}^{\text{v}}$. In contrast, active viewpoint movement introduces an additional influence from the viewpoint itself, resulting in the mapping $f^{\text{active}}:\mathcal{S}\times\mathcal{V}\rightarrow\mathcal{O}^{\text{v}}$. Since the 3D-CNN lacks spatial rotation invariance, voxelizing the scene using the world coordinate system as the reference frame can cause features of similar structures from different viewpoints to appear distant in the feature space after encoding, as shown in Fig.4. This misalignment negatively impacts the overall training efficiency of the model.
@@ IV-C Viewpoint-aware Demo Augmentation
To provide early guidance and reduce the need for extensive exploration, we manually alternate moving the camera and gripper to generate and collect a small number of demonstration trajectories as high-quality examples. Each demo trajectory $\tau_{\text{d}}=\{(\mathbf{o}_{i},\mathbf{a}_{i})\}^{N}$ consists of “observation-action” pairs collected at a fixed sampling rate. Existing demonstration augmentation methods[1][2][31] typically involve two main processes: key frame discovery and observation sampling. In the key frame discovery process, a limited set of key frame indices is selected from a trajectory $\tau_{d}$, and transition samples are constructed based on the corresponding “observation-action” pairs. The observation sampling process then generates additional transition samples by resampling the initial observations between the identified key frames. However, existing methods do not account for viewpoint movement. Our method incorporates key frame discovery and observation resampling, taking into account viewpoint movement, enabling viewpoint-aware demonstration augmentation. Specifically, we define the key frame discovery function as $KF:\tau_{\text{d}}\rightarrow\big{\{}(k_{j}^{\text{c}},k_{j}^{\text{g}})\big{\}}^{M}\subset\mathbb{R}_{+}^{2}$, which retrieves the key frames of both viewpoint and gripper movements in the demonstration trajectory based on predefined rules.
The discovery of gripper key frames $k_{j}^{\text{g}}$ follows the method described in [1]. We define key viewpoints as the first stable viewpoint between every two gripper key frames, corresponding to the frame index $k_{j}^{\text{c}}$. The relationship between key frame indices satisfies $10\leq k_{j-1}^{\text{c}}\leq k_{j-1}^{\text{g}}\leq k_{j}^{\text{c}}\leq k_{j}^{\text{g}}\leq N-$ 1. As shown in Fig.5, after the key frame discovery process, we construct the raw transition samples $\text{T}_{j}$ formalized in Eq.(5).
$\begin{split}\text{T}_{j}&=\{\text{T}_{j}^{\text{v}},\text{T}_{j}^{\text{p}},r_{j}\}\\ &=\{<\mathbf{o}_{k_{j-1}^{\text{g}}},\mathbf{a}_{k_{j-1}^{\text{g}}}^{\text{c}},\mathbf{o}_{k_{j}^{\text{c}}}>,<\mathbf{o}_{k_{j}^{\text{c}}},\mathbf{a}_{k_{j}^{\text{c}}}^{\text{g}},\mathbf{o}_{k_{j}^{\text{g}}}>,r_{j}\}\end{split}$ (5)
To generate additional augmented transitions, we uniformly sample within the intervals $k_{j-1}^{\text{g}}\sim k_{j}^{\text{c}}$ and $k_{j}^{\text{c}}\sim k_{j}^{\text{g}}$, obtaining the augmented frame indices$(k_{j-1}^{\text{g}})^{\text{aug}}$ and $(k_{j}^{\text{c}})^{\text{aug}}$. The resulting augmented transition $\text{T}_{j}^{\text{aug}}$ is formalized in Eq.(6).
$\begin{split}\text{T}_{j}^{\text{aug}}&=\{(\text{T}_{j}^{\text{v}})^{\text{aug}},(\text{T}_{j}^{\text{p}})^{\text{aug}},r_{j}\}\\ &=\{<\mathbf{o}_{(k_{j-1}^{\text{g}})^{\text{aug}}},\mathbf{a}_{(k_{j-1}^{\text{g}})^{\text{aug}}}^{\text{c}},\mathbf{o}_{(k_{j}^{\text{c}})^{\text{aug}}}>,\\ &<\mathbf{o}_{(k_{j}^{\text{c}})^{\text{aug}}},\mathbf{a}^{\text{g}}_{(k_{j}^{\text{c}})^{\text{aug}}},\mathbf{o}_{k_{j}^{\text{g}}}>,r_{j}\}\end{split}$ (6)
Figure 6: Viewpoint adjustment and ROI selection in a 2D simplified schematic scene. We encourage the model to select non-empty ROIs (green dashed boxes) and maximize the number of occluded voxels within the ROI that can be observed from the adjusted viewpoint $\mathbf{v}_{t^{\prime}}$ compared to the original viewpoint $\mathbf{v}_{t}$. This strategy promotes active interaction with the environment and enables the model to capture more structural information within the ROI.
@@ IV-D Task-agnostic Auxiliary Reward
We use the task completion signal as a sparse task reward to avoid designing targeted rewards for different tasks; however, sparse reward only provides positive feedback to the model at the end of an episode, making it difficult to efficiently guide the model in sensor-motor joint learning. To address this, we introduce two task-agnostic auxiliary rewards that provide real-time feedbacks to the training process.
To encourage effective interaction with the environment and accurate environment modeling, we propose the Interaction Encouragement Reward, denoted as $r_{\text{i}}$ As illustrated in Fig.6, reward $r_{\text{i}}$ provides a positive signal when the NBV policy infers an ROI in a reachable and non-empty area of the scene, and penalizes actions that select an ROI in an empty or unreachable area due to distance or trajectory planning failures. The specific value assignment for $r_{\text{i}}$ is detailed in Eq.(7).
$r_{\text{i}}=\begin{cases}0.02,&\text{ROI reachable and non-empty}\\ 0.0,&\text{ROI reachable but empty}\\ -0.02,&\text{ROI not reachable and empty}\end{cases}$ (7)
To encourage the model to select viewpoints that capture more structural information from the ROI, we introduce the Entropy Reduction Reward , denoted as $r_{\text{e}}$, inspired by [24]. The core idea is to reduce the occupancy entropy within the ROI after adjusting the viewpoint, as illustrated in Fig.6. Specifically, assuming the ROI contains $N$ voxels, the occupancy entropy of the $i$-th voxel is calculated using the Shannon entropy function, as shown in Eq.(8), where $p_{\text{o}}(i)$ represents the voxel’s occupancy probability. If a voxel is occluded, its occupancy probability is set to $p_{\text{o}}=0.5$.
@@
We define the reward $r_{\text{e}}$ as Eq.(10), where the coefficient $k$ is empirically set to 1.0.
$r_{\text{e}}=k\cdot\Delta E,k=1.0$ (10)
The auxiliary reward $r_{\text{i}}$ necessitates the cooperation of both the NBV and NBP policies, whereas $r_{\text{e}}$ is exclusive to the NBV policy. Therefore, the reward for updating the NBP policy is expressed as $r_{\text{nbp}}=r_{\text{task}}+r_{\text{i}}$, while the reward for updating the NBV agent is $r_{\text{nbv}}=r_{\text{task}}+r_{\text{e}}+r_{\text{i}}$, where $r_{\text{task}}\in[0,1]$ is determined by task completion. It is important to note that the introduced rewards are designed to be task-agnostic, which ensures that the rewards can generalize across different tasks without the need for manual tuning.
Figure 7: Reletive task occlusion rate of selected tasks. We use the active viewpoint occlusion rate of 30 demonstration trajectories for each task as a benchmark. Based on this, we calculate the relative task occlusion rates for the front, overhead, and wrist viewpoints in the demonstration trajectories of each task.
@@ V Experiment Setup
V-A Task Setting
We trained and evaluated our model on RLBench [9], a public benchmark designed for reinforcement learning manipulation algorithms. When selecting evaluation tasks, we considered two primary criteria: 1) Single task: each task must have at least one fixed viewpoint with a high level of visual occlusion. 2) Task range: no single fixed or wrist-mounted camera can provide a low occlusion rate for all tasks. Since the region of interest (ROI) the robot focuses on may change during different manipulation phases, and there may be no clear manipulation objects (e.g., when the gripper is transferring in midair), it is challenging to use an image segmentor to locate objects and define the task occlusion rate as in previous methods[14][24]. To address this, we define the Task Occlusion Rate (TOR) for a given viewpoint as the average occupancy entropy of the ROI where the gripper’s goal actions occur, based on observations from that viewpoint across $L=100$ demo trajectories, as described in Eq.(11), Here, $E^{l}_{t^{\prime}_{k}}$ represents the ROI occupancy entropy at the $k$-th keyframe in trajectory $l$ after viewpoint adjustment where the ROI center is set to the gripper’s goal position in the demo observation. To compare the occlusion levels across different tasks, we assume that the collected demo trajectories contain the strongest task-specific prior knowledge. Thus, we can obtain the Relative TOR by dividing the TOR of the selected viewpoint by the TOR of the human-demoed active viewpoint (oracle).
@@ V-A Task Setting
Fig.7 presents the relative TOR statistics for 8 selected RLbench tasks under common overhead, front , wrist-mounted cameras as well as the human-demoed active camera, based on $100$ demo trajectories. The results show that, due to task characteristics and initial scene randomization, both fixed and wrist viewpoints experience varying degrees of occlusion across the 8 tasks. In contrast, the active camera demonstrates the potential to achieve lower task occlusion.
V-B Baselines
We compaired our model against two categories baseline models: 1) visual manipulaiton model that utilze common used static third-person cameras and first-person wrist-mounted camera; 2) Active third-person camera based manipulation models.
For the first category, we compared our active vision-action model against the recent visual-based C2FARM [2] model with static front, overhead camera as well as the wrist-mounted camera settings. This comparison aims to explore the potential effectiveness of our active vision-action model in handling tasks with limited observability.
For the second category, using the same third-person active camera setting, we firstly compared our model against the Behavior Cloning (BC) model. Through this comparison, we seek to investigate whether our model not only mimics the given limited demonstrations but also learns new sensor-motor skills through interaction with the environment. Additionally, we tested a Random viewpoint model, following [3], which shares the same architecture as ours but with the learning of the NBV policy network disabled (learning rate set to 0). This setup allows us to evaluate the importance of the active viewpoint selection in the NBV policy and to assess how much performance improvement is achieved through active viewpoint adjustment as opposed to random viewpoint choices.
V-C Training Setting
We first collected 30 demo trajectories for each task and stored them in the replay buffer as prior samples or the training dataset, and all models were trained from scratch.
The implementation of C2FARM was based on their official release and trained using sparse task rewards. The Behavior Cloning model was trained with the Mean Squared Error (MSE) loss, while the Random model utilized the same reward structure as our model, as described in SectionIV-D. All models were trained using the Adam optimizer with an initial learning rate of $0.0005$, with no learning rate decay applied during training. Each model was trained for 40,000 updates with a batch size of 128, and all experiments were conducted on a machine equipped with 4 NVIDIA GTX 4090 GPUs.
TABLE I: Task success rate: Our model outperform baselines on most tasks and on average . Mean and std of 100 tests are reported.
@@ V-C Training Setting
TABLE II: Task episode length. The maximum length of each task episode is limited to 10 steps. Mean and std of 100 tests are reported.
@@ VI Result
Main Performance. We first analyzed the average task success rate and the average episode length across different algorithms, which provide a macro-level assessment of the models’ proficiency in learning sensor-motor coordination.
Table I presents the average success rate for tasks across 100 testing episodes under 3 evaluation seeds. Our model achieved the highest average success rate across the eight tasks, outperforming the second-best by 21%. It also exhibited relatively high manipulation performance in scenarios with significant occlusion from fixed viewpoints, such as meat_on_grill, meat_off_grill and unplug_charger.
Similarly, Table II shows that the average episode length for tasks follows the same trend as the success rate. Our model achieved the shortest average manipulation steps across all tasks, outperforming the second-best model by 28.2%. Furthermore, it demonstrated the lowest variance in manipulation steps, highlighting the model’s high proficiency and consistency after training.
Moreover, we observed that in tasks where the occlusion rate is low or the scene object positions are relatively predictable, fixed and wrist-mounted cameras benefit from smaller variance in observation samples, leading to enhanced skill acquisition. For example, in tasks such as basketball_in_hoop, press_switch, put_rubbish_in_bin, the C2FARM model achieved comparable success rates from both front and overhead views, occasionally even surpassing our model. However, it is important to note that when these fixed cameras fail to capture adequate observations, performance tends to decline sharply, leading to higher performance variance across the 8 tasks.
In contrast, our model is less sensitive to visual occlusion and consistently maintains strong manipulation performance. Under the active viewpoint setting, although the BC model demonstrated some ability to adjust its viewpoint (as detailed below), its effectiveness was constrained by the limited number of training samples, which restricted its capacity to learn interaction skills. As a result, it frequently failed to complete the assigned tasks, achieving an average success rate of only 14%. Meanwhile, the random viewpoint model performed poorly across all 8 tasks, exhibiting the weakest manipulation performance, indicating its inability to effectively learn sensor-motor coordination skills.
Figure 8: Viewpoint distribution overview: Red circles indicate the distribution of manually set viewpoints across 30 demo trajectories, while blue triangles represent the distribution of viewpoint positions during successful episodes of our model over 100 test episodes. The model’s inferred viewpoints do not strictly mimic the demonstration trajectories but instead exhibit a degree of refinement and generalization.
@@
Figure 9: Partial rollout cases of our model. The figure presents the image observations corresponding to the first three inferred viewpoints $\mathbf{v}^{*}$ and the associated gripper actions $\mathbf{g}^{*}$ for four representative tasks during the test episodes.
TABLE III: A comparison of information gain to baselines.
      Models              ASSIG       (avg/max)              AIIG       (avg/max)
      C2FARM(wrist)       0.0/0.0       N/A
      C2FARM(overhead)       0.0/0.0       0.0/0.0
      C2FARM(front)       0.0/0.0       0.0/0.0
      Random(active)       0.0/5.1%       0.0/5.2%
      BC(active)       0.7%/2.3%       5.6%/18.1%
      Ours(active)       3.5%/6.9%       6.8%/25.1%
@@
Active Perception Strategy Analysis. We further analyzed the perception strategies of all models during the operation process using two metrics: Average Single Step Information Gain (ASSIG) and Average Initial Information Gain (AIIG). These metrics measure the models’ ability to acquire observation information through multi-step and single-step interactions, respectively. Specifically, we define ASSIG as the average reduction in entropy within the ROI before and after a single viewpoint movement.Similarly, the initial information gain is defined as the average reduction in entropy of the ROI after a viewpoint movement, compared to the initial viewpoint. These two metrics are given by:
@@
Where $L=100$ denotes the number of testing episodes, $K_{l}$ represents the number of steps required to complete episode $l$, and $E^{l}_{t_{k}}$ is defined in Eq.(9). The term $[E^{l}_{t_{k}}]_{\mathbf{v}{t_{0}}}$ represents the visual entropy of the current ROI from the episode’s initial viewpoint. The statistical results in Table III show that our model significantly improves the observation of the ROI region by independently adjusting camera viewpoints, outperforming the second-best BC model by 2.8% and 1.2% in the ASSIG and AIIG scores, respectively. In contrast, static front and overhead cameras, which are unable to move, fail to achieve either single-step or initial information gain through viewpoint shifts. Similarly, the wrist-mounted camera, being unable to move independently of the gripper, is limited to inferring the next gripper actions based on the current viewpoint. This results in a zero ASSIG score and makes it unsuitable for AIIG computation.
Fig.8 illustrates the distribution of viewpoints learned by our model alongside the key viewpoints distribution from 30 demo trajectories. Notably, the viewpoints learned by our model do not strictly follow the patterns observed in the demonstration trajectories. For tasks such as put_rubbish_in_bin, press_switch, unplug_charger, and meat_off_grill, the learned viewpoints are more concentrated. In contrast, for tasks like meat_on_grill, open_drawer, and take_money_out_safe, the model explores a broader range of viewpoints. This variation indicates that our model has effectively learned to select task-relevant viewpoints through interactive learning, rather than merely replicating the active perception patterns from the demonstration data. Furthermore, Fig. 8 presents camera observations at keyframe moments of partial task episodes, showcasing our model’s capacity for active visual perception during task execution.
Figure 10: Comparison of success, timeout, and execution failure rate across models.
@@
Environmental Modeling and Interactive Capabilities. To further evaluate the environment modeling and interaction capabilities of various models, we recorded the success rate, timeout rate, and execution failure rate over 100 test episodes for each model, as illustrated in Fig.10. The C2FARM model showed a high likelihood (up to 42%) of episodes terminating prematurely due to failures in low-level trajectory execution. Similarly, the random viewpoint model saw the majority of its test episodes end in failure because of execution errors. While execution failures accounted for only 4% of the BC model’s episodes, as many as 82% of the episodes failed due to timeouts. However, our model showed only a 13% rate of episodes ending prematurely due to trajectory execution failures and a mere 9% failure rate due to timeouts. These results suggest that our model effectively integrates low-level motion execution considerations when inferring manipulation actions, resulting in more robust performance across the test episodes.
Additionally, as shown in Table IV, we analyzed the average steps to first interaction with scene objects and the average number of non-interactive steps during testing for each model. Our findings indicate that the behavior cloning exhibited a relatively high number of non-interactive steps, which aligns with the high timeout rate shown in Fig.10. This pattern suggests that behavior cloning struggles to learn precise interactive manipulations from a limited number of demonstration trajectories, leading to frequent non-interactive steps and episode timeouts. In contrast, our model consistently initiates interaction with the environment more quickly and significantly reduces the time spent on non-interactive actions. These results demonstrate that, after training, our model achieves superior accuracy in modeling environmental dynamics and understanding structural relationships.
TABLE IV: First interaction step and non-interaction step comparison.
Models First Interaction Step Non-interaction Step
C2FARM(wrist) 3.21±1.82 4.86±2.70
C2FARM(overhead) 2.97±1.70 4.41±2.77
C2FARM(front) 3.41±2.28 4.17±2.74
Random(active) 9.07±1.47 9.49±1.07
BC(active) 6.31±2.46 7.54±2.15
Ours(active) 1.64±0.61 2.15±1.03
@@
Ablation Study In this section, we investigate the impact of three key strategies on manipulation performance and various evaluation metrics: (1) viewpoint-centric voxel alignment (Section IV-B), (2) viewpoint-aware demon augmentation (Section IV-C), and (3) task-agnostic auxiliary rewards (Section IV-D). To assess the individual and combined contributions of these strategies, we conducted separate ablation tests for each, followed by a comprehensive test removing all three strategies.
As shown in Table V, each of the three strategies contributed to notable improvements in the success rate of our proposed model. Among the three, voxel alignment and viewpoint-aware demonstration augmentation had the most substantial impact, with the success rate dropping by approximately 40% when either of these strategies was ablated, highlighting their critical roles. The auxiliary reward strategy, while less significant than the other two, still provided approximately a 6% improvement, demonstrating its contribution to overall performance.
@@
Beyond success rate improvements, all three strategies contributed to reductions in episode length, overtime ratio, planning failure rate, and the average number of non-interactive steps per episode. This demonstrates their broader importance to the algorithm’s effectiveness. Interestingly, while the voxel alignment strategy significantly boosts overall performance, it leads to a reduction in ASSIG score. We hypothesize that without viewpoint alignment, the NBV agent avoids potential issues with out-of-bounds actions and clipping, which may simplify the learning of viewpoint transformations. However, the overall performance still confirms that viewpoint alignment plays a crucial role in improving operational effectiveness.
When all three strategies were ablated, the performance across the eight tasks showed the most severe decline, with the average success rate dropping to just 22%, even lower than the static-viewpoint baseline. This highlights the synergistic benefits of these three strategies, as their combination significantly boosts the model’s overall performance.

### v2

@@ IV-C Viewpoint-aware Demo Augmentation
To provide early guidance and reduce exploration, we manually alternate camera and gripper movements to generate a small set of demonstration trajectories. Each demo trajectory $\tau_{\text{d}}=\{(\mathbf{o}_{i},\mathbf{a}_{i})\}^{N}$ consists of observation-action pairs collected at a fixed rate. To further leverage limited demo data, existing augmentation methods [1, 2, 28] first select key frame indices from the demo trajectory, construct transition samples from the key observation-action pairs, and then resample between key frames to generate additional transitions. However, they cannot handle the observation variance caused by viewpoint shifts. We define the key frame discovery function as $KF:\tau_{\text{d}}\rightarrow\big{\{}(k_{j}^{\text{c}},k_{j}^{\text{g}})\big{\}}^{M}\subset\mathbb{R}_{+}^{2}$, identifying key frames for both viewpoint and gripper movements. Gripper key frames $k_{j}^{\text{g}}$ are marked when the gripper is stationary or its closure state changes, while key viewpoints are defined as the first stable viewpoint between two gripper key frames, indexed by $k_{j}^{\text{c}}$. The key frame indices satisfy $10\leq k_{j-1}^{\text{c}}\leq k_{j-1}^{\text{g}}\leq k_{j}^{\text{c}}\leq k_{j}^{\text{g}}\leq N-$ 1. As shown in Fig.5, raw transition samples $\text{T}_{j}$ are then constructed as formalized in Eq.(5).
$\begin{split}\text{T}_{j}&=\{\text{T}_{j}^{\text{v}},\text{T}_{j}^{\text{p}},r_{j}\}\\ &=\{(\mathbf{o}_{k_{j-1}^{\text{g}}},\mathbf{a}_{k_{j-1}^{\text{g}}}^{\text{c}},\mathbf{o}_{k_{j}^{\text{c}}}),(\mathbf{o}_{k_{j}^{\text{c}}},\mathbf{a}_{k_{j}^{\text{c}}}^{\text{g}},\mathbf{o}_{k_{j}^{\text{g}}}),r_{j}\}\end{split}$ (5)
To generate additional augmented transitions, we uniformly sample within the intervals $k_{j-1}^{\text{g}}\sim k_{j}^{\text{c}}$ and $k_{j}^{\text{c}}\sim k_{j}^{\text{g}}$, obtaining the augmented indices$(k_{j-1}^{\text{g}})^{\text{aug}}$ and $(k_{j}^{\text{c}})^{\text{aug}}$. The resulting augmented transition $\text{T}_{j}^{\text{aug}}$ is formalized in Eq.(6).
$\begin{split}\text{T}_{j}^{\text{aug}}&=\{(\text{T}_{j}^{\text{v}})^{\text{aug}},(\text{T}_{j}^{\text{p}})^{\text{aug}},r_{j}\}\\ \end{split}$ (6)
@@ IV-D Task-agnostic Auxiliary Reward
We use sparse reward to avoid designing task-specific rewards. However, since sparse reward only provides feedback at the end of an episode, it hinders efficient sensor-motor learning. Here we introduce two task-agnostic auxiliary rewards to provide real-time guidance to the training process.
To encourage interaction with the environment and accurate environment modeling, we propose the Interaction Encouragement Reward, denoted as $r_{\text{i}}$ As illustrated in Fig.6, $r_{\text{i}}$ provides positive signals when the NBV policy selects a reachable, non-empty ROI and penalizes actions targeting empty or unreachable areas due to execution failures. The specific value assignment for $r_{\text{i}}$ is detailed in Eq.(7).
$r_{\text{i}}=\begin{cases}0.02,&\text{ROI reachable and non-empty}\\ 0.0,&\text{ROI reachable but empty}\\ -0.02,&\text{ROI unreachable}\end{cases}$ (7)
To encourage the model to select viewpoints that capture more structural information from the ROI, we introduce the Entropy Reduction Reward($r_{\text{e}}$), inspired by [21]. The goal is to reduce occupancy entropy within the ROI after viewpoint adjustment, as shown in Fig.6. Assuming the ROI contains $N$ voxels, the occupancy entropy of the $i$-th voxel($e_{i}$) is calculated using the Shannon entropy function(Eq.(8)), where $p_{\text{o}}(i)$ represents its occupancy probability. If a voxel is occluded, $p_{\text{o}}$ is set to 0.5.
@@
We define reward $r_{\text{e}}=k\cdot\Delta E$, where $k$ is empirically set to 1.0. The auxiliary reward $r_{\text{i}}$ requires cooperation between the NBV and NBP policies, while $r_{\text{e}}$ is exclusive to the NBV policy. Therefore, the reward for updating the NBP policy is expressed as $r_{\text{nbp}}=r_{\text{task}}+r_{\text{i}}$, and the NBV reward is $r_{\text{nbv}}=r_{\text{task}}+r_{\text{e}}+r_{\text{i}}$, where $r_{\text{task}}\in[0,1]$ is determined by task completion. These task-agnostic rewards generalize across tasks without the need for manual tuning.
Figure 6: Viewpoint adjustment and ROI selection in a 2D simplified schematic scene. The model is encouraged to select non-empty ROIs and reduce occluded voxels from the adjusted viewpoint $\mathbf{v}_{t^{\prime}}$.
@@ V Experiment Setup
Task Setting. We trained and evaluated our model on RLBench[9], a widely-used public benchmark for robotic manipulation in reinforcement learning. Task selection followed two criteria: 1) each task must include at least one viewpoint with significant visual occlusion, and 2) no single fixed or wrist-mounted camera provides consistently low occlusion across selected tasks. As the ROI shifts throughout the manipulation process, and there may be phases without distinct objects(e.g., midair gripper movements), defining occlusion rates based on object segmentation, as done in previous methods[14, 21], is impractical. To address this, we define Task Occlusion Rate(TOR) for a given viewpoint as the average occupancy entropy of the ROI where the gripper’s goal actions occur, based on observations from that viewpoint across $L=100$ demo trajectories, as described in Eq.(10). Here, $E^{l}_{t^{\prime}_{k}}$ represents the ROI occupancy entropy at the $k$-th keyframe in trajectory $l$ after viewpoint adjustment where the ROI center is set to the gripper’s goal position in the demo observation. To compare the occlusion levels across selected tasks, we assume that the human-demoed active viewpoint contains the strongest task-specific prior knowledge. Thus, we can obtain the Relative TOR by dividing the TOR of the selected viewpoint by the TOR of the human-demoed active viewpoint(oracle).
@@
Fig.7 presents the relative TOR statistics for 8 selected RLbench tasks under overhead, front, wrist-mounted cameras as well as the human-demoed active camera.
Figure 7: Reletive task occlusion rate of selected tasks. We use the active viewpoint occlusion rate of 100 demonstration trajectories for each task as a benchmark.
Baselines. We compared our model against two baseline categories: 1) visual manipulation model that utilize static third-person and wrist-mounted cameras, and 2) active third-person camera-based models. For the first category, we compared our model against C2FARM [2] with static front, overhead and wrist-mounted camera settings to explore the potential effectiveness of our active vision-action model in handling tasks with limited observability. For the second category, we compared our model with Behavior Cloning(BC) to test whether it can learn new sensor-motor skills beyond demonstration imitation. We also tested a Random viewpoint model [3], which shares the same architecture as ours but with the NBV policy network disabled(learning rate set to zero). This setup allows us to assess how much performance improvement is achieved through active viewpoint adjustment as opposed to random choices.
Training Setting. We collected 30 demo trajectories for each task and used them as prior knowledge for training. C2FARM was implemented based on its official release and trained with sparse task rewards. The Behavior Cloning model was trained with MSE loss, and the Random model used the same reward structure as our model(SectionIV-D). All models were trained using Adam with a learning rate of $0.0005$ for 40,000 updates, with a batch size of 128. Experiments were conducted on a machine with 4 NVIDIA RTX 4090 GPUs.
TABLE I: Task success $\uparrow$ . Mean and std of 100 tests are reported(3 seeds).
@@
TABLE II: Task episode $\downarrow$ . The maximum length of each task episode is limited to 10 steps.
@@ VI Result
Main Performance. We first analyzed the average task success rate(SR) and episode length(EL) across different models to evaluate overall sensor-motor performance. Table I shows that our model achieved the highest success rate, outperforming the second-best by 21%, especially in occluded tasks like meat_on_grill, meat_off_grill and unplug_charger. Similarly, Table II shows that our model had the shortest average manipulation steps, outperforming the second-best by 28.2% and showing the least variance in step counts.
Fixed and wrist-mounted cameras performed well in low-occlusion tasks, such as basketball_in_hoop, press_switch, put_rubbish_in_bin, C2FARM occasionally matching or surpassing our model. However, their performance dropped sharply in occluded scenarios, leading to a higher variance. In contrast, our model remained robust in occluded tasks. Although the BC model demonstrated some ability to adjust its viewpoint(as detailed below), it struggled with limited training samples, achieving only a 14% success rate. The random viewpoint model performed the worst across all tasks, highlighting its inability to learn sensor-motor coordination.
@@
Figure 8: Partial rollout cases and viewpoint distribution. (a) The image observations corresponding to the first three inferred viewpoints $\mathbf{v}^{*}$ and the gripper actions $\mathbf{g}^{*}$ for four representative tasks during the test episodes. (b) Red circles indicate manually set viewpoints from 30 demo trajectories, while blue triangles represent the learned viewpoint positions from successful episodes across 100 test runs.
@@
Active Perception Strategy Analysis. We further analyzed the perception strategies of all models using two metrics: Average Single Step Information Gain(ASSIG) and Average Initial Information Gain(AIIG). ASSIG captures the average reduction in ROI entropy before and after a single viewpoint shift, while AIIG measures the entropy reduction compared to the initial viewpoint.
@@
Where $L=100$ denotes the number of testing episodes, $K_{l}$ represents the length of episode $l$, and $E^{l}_{t_{k}}$ is defined in Eq.(9). The term $[E^{l}_{t_{k}}]_{\mathbf{v}{t_{0}}}$ represents the visual entropy of the current ROI from the episode’s initial viewpoint. As shown in Table III, our model significantly enhances ROI observation by adjusting viewpoints, outperforming the BC model by 2.8% in ASSIG and 1.2% in AIIG. In contrast, static front and overhead cameras, which cannot shift viewpoints, fail to achieve information gain. The wrist-mounted camera, unable to move independently of the gripper, shows zero ASSIG and is unsuitable for AIIG computation.
Fig.8(b) compares the distribution of viewpoints learned by our model with the manually set active viewpoints from the demo trajectories used during training. Unlike demo patterns, our model’s learned viewpoints are more concentrated in some tasks(Basketball_in_hoop, unplug_charger) and more varied in others(meat_on_grill, open_drawer), indicating it learned to select task-relevant viewpoints rather than simply imitating demo data. The camera observations at key moments(Fig. 8) further showcase our model’s active visual perception capabilities.
Figure 9: Proportion of success, timeouts, and execution failures across test episodes for different models.
TABLE III: Comparison of information gain and interaction capacity across different models.
Models ASSIG $\uparrow$ (avg/max) AIIG $\uparrow$ (avg/max) First $\downarrow$ Interaction Non- $\downarrow$ interaction
C2FARM(wrist) 0.0/0.0 N/A 3.21±1.82 4.86±2.70
C2FARM(overhead) 0.0/0.0 0.0/0.0 2.97±1.70 4.41±2.77
C2FARM(front) 0.0/0.0 0.0/0.0 3.41±2.28 4.17±2.74
Random(active) 0.0/5.1% 0.0/5.2% 9.07±1.47 9.49±1.07
BC(active) 0.7%/2.3% 5.6%/18.1% 6.31±2.46 7.54±2.15
Ours(active) 3.5%/6.9% 6.8%/25.1% 1.64±0.61 2.15±1.03
@@
Environmental Modeling and Interactive Capabilities. We evaluated the models’ environment modeling and interaction capabilities by recording success, timeout(TO), and execution failure(EF) rates over 100 test episodes(Fig.9). The C2FARM model exhibited a high failure rate(up to 42%) due to trajectory execution failures. Similarly, the random viewpoint model saw the majority of its test episodes end in failure because of execution errors. Although only 4% of the BC model’s episodes ended due to execution errors, 82% failed due to timeouts. In contrast, our model showed a 13% failure rate from trajectory errors and only 9% from timeouts, indicating effective integration of low-level motion execution.
As shown in Table III, we also analyzed the average steps to first interaction(FI) and non-interactive steps(NI) per model. The BC model exhibited frequent non-interactive steps, correlating with its high timeout rate, suggesting difficulty learning precise manipulations from limited demonstrations. In contrast, our model initiated interactions more quickly, reducing non-interactive steps, and demonstrating superior accuracy in modeling environmental dynamics and structural understanding after training.
@@
Ablation Study. We evaluated the impact of three strategies on manipulation performance:(1) viewpoint-centric voxel alignment(Section IV-B), (2) viewpoint-aware demo augmentation(Section IV-C), and (3) task-agnostic auxiliary rewards(Section IV-D). Separate ablation tests were conducted for each, as well as a combined test removing all strategies.
Table IV shows that each strategy effectively improved success rates(SR), with voxel alignment and viewpoint-aware augmentation contributing the most, as success rates dropped by approximately 40% without them and auxiliary rewards added around 6% improvement.
@@
Beyond success rate improvements, all three strategies contributed to reductions in episode length(EL), timeouts rate(TO), execution failures(EF), and the average number of non-interactive steps(NI) per episode. This demonstrates their broader importance to the model’s effectiveness. Interestingly, while the voxel alignment strategy significantly boosts overall performance, we found it leads to a reduction in ASSIG score. We hypothesize that without viewpoint alignment, the NBV agent avoids potential issues with out-of-bounds actions and clipping, which may simplify the learning of viewpoint transformations. However, the overall performance still confirms that viewpoint alignment plays a crucial role in improving operational effectiveness.
Ablating all strategies led to a sharp decline, reducing the average success rate to 22%, underscoring the synergistic effect of combining these strategies for optimal performance.